OpenStudy (anonymous):

Find the length of the radius of the following circle. x² + y² = 16 4 8 16

OpenStudy (anonymous):

Standard form of a circle: \[(x-h)^{2}+(y-k)^{2}=r^{2}\] Your circle: \[(x-0)^{2}+(y-0)^{2}=(\sqrt{16})^{2}\] So: \[r=\sqrt{16}=4\]